Please help me solve this equation and explain how you did it: 9x^2 + 6x +1 =64
Answer by Alan3354(69443)

9x^2 + 6x +1 =64
9x^2 + 6x - 63 = 0
3x^2 + 2x - 21 = 0
(x + 3)*(3x - 7) = 0
x = -3
x = 7/3